### v2

# Extending the SAREF4ENER Ontology with Flexibility Based on FlexOffers

Abstract
A key element to support the increased amounts of renewable energy in the energy system is flexibility, i.e., the possibility of changing energy loads in time and amount. Many flexibility models have been designed; however, exact models fail to scale for long time horizons or many devices. Because of this, the FlexOffers model has been designed, to provide device-independent approximations of flexibility with good accuracy, and much better scaling for long time horizons and many devices. An important aspect of the real-life implementation of energy flexibility is enabling flexible data exchange with many smart energy appliances and market systems, e.g., in smart buildings. For this, ontologies standardizing data formats are required. However, the current industry standard ontology for integrating smart devices for energy purposes, SAREF for Energy Flexibility (SAREF4ENER), only has limited support for flexibility and thus cannot support important use cases. In this paper, we propose an extension of SAREF4ENER that integrates full support for the complete FlexOffer model, including advanced use cases, while maintaining backward compatibility. This novel ontology module can accurately describe flexibility for advanced devices such as electric vehicles, batteries, and heat pumps. It can also capture the inherent uncertainty associated with many flexible load types.
I Introduction
The last decades have seen an increasing usage of renewable energy, and the development of new technologies and concepts to enable its exploitation. In particular, among those new concepts, one has crucial importance for handling the intermittency that comes with renewable energy generation: flexibility [1], i.e., the capability of changing energy loads in time or amount of energy. Flexibility enables optimizing energy consumption to achieve several objectives: renewable energy usage maximization [2], peak shaving [3, 4], CO2 emissions minimization [5], energy cost minimization [6, 7], provision of ancillary services [8], congestion prevention [9], and demand response [10]. Several models have been developed for flexibility. Many of them are accurate in their description of flexibility, but scale poorly with respect to the length of time horizons and the number of devices. A model that solves these issues is FlexOffers [11], a device-independent model that represents flexibility by using energy constraints that approximate the amount of available flexibility, but scales very well in scenarios with many devices or long time horizons. FlexOffers generates flexoffers (FOs) objects, which describe flexibility from the considered devices.FlexOffers is a powerful model used in more than 30 European and national projects, whose real-life implementation has been advancing in the last decade. An important aspect is its interaction with smart devices/buildings and the ontologies representing them. To date, a widely used ontology for smart buildings is the Smart Application REFerence (SAREF) [12] ontology. Specifically, several extensions of SAREF have been created for different domains: SAREF for Energy (SAREF4ENER) [13] has been designed to represent energy devices and services, and is the de facto standard for that. However, SAREF4ENER fails to capture energy flexibility for many use cases, specifically when flexibility constraints become more complex: this is the case for advanced devices such as batteries, electric vehicles (EVs), and heat pumps. This raises the need for an ontology capable of handling these cases. To meet this challenge, an ontology called domOS Common Ontology (dCO [^1] https://w3id.org/dco[14] has been proposed to ensure full flexibility support in the interoperability of smart buildings and services. The contribution we bring in this paper is the dCO flexibility module: this module extends SAREF4ENER in its representation of flexibility services, enabling compatibility with the full Flexoffers model and constraints, and ensuring that the most relevant flexible devices and use cases for them are represented. In this paper, we show how the flexibility module of dCO extends SAREF4ENER, how it represents the full Flexoffer model and constraints, and some advanced use cases that can be represented by it. The remainder of this paper is organized as follows. Section II describes FOs and the SAREF4ENER and dCO ontologies, while Section III shows how dCO extends SAREF4ENER and enables full FO representation. Section IV evaluates the dCO flexibility module, showing which use cases can and cannot be represented by FOs and SAREF4ENER. Finally, Section V concludes the work and provides ideas for future research.
II Preliminaries
II-A Running Example
We start by making a running example so that we can describe the Flexoffer concept in the following subsections. Here, we will consider the case of a heat pump operating in a room. The heat pump is a Daikin Altherma [15], with maximum operational power $P_{max}=4.6$ kW and coefficient of performance (COP) $3.65$. This model supports the SG-Ready interface [16, 17], a standardized smart grid communication protocol that enables seamless integration with modern energy management systems. The room has a rectangular section with sides of $4$ and $5$ meters respectively, and it is $3$ meters high. The heat transfer coefficient is $c_{ht}=6\frac{W}{m^{2}K}$ [18], and heat is dispersed through one of the $3$m x $4$m side walls. The room temperature is $295$K ($22^{\circ}$), while the outside temperature is $275$K ($2^{\circ}$). We assume that room temperature must stay between $293$K ($20^{\circ}$) and $297$K ($24^{\circ}$), respectively, because of user settings. For reference, the amount of energy (heat) needed to maintain the temperature to $295$K for one hour can be obtained by
$H=(3\cdot 4)m^{2}\cdot 6\frac{W}{m^{2}K}\cdot(295K-275K)\cdot 3600s=1.44kWh.$
Since $COP=3.65$, the amount of electrical energy needed to provide this amount of heat is $E=\frac{H}{COP}=0.395kWh.$
II-B FlexOffer Life Cycle
The Flexoffers model describes flexibility in a unified format for different device types [11, 19], by providing an approximated representation of flexibility. In exchange for the small amount of accuracy lost with respect to exact models, modeling flexibility with FOs allows for much better scalability over long time horizons and high numbers of devices. FOs provide constraints for energy consumption: optimizing an FO means solving an optimization problem on energy consumption (e.g., cost minimization) within the constraints given by the FO. The solution of this optimization problem is a set of values $(e_{1},\ldots,e_{T})$ for energy consumption at each time unit: this is called a schedule. FOs can also be aggregated: this means that, from $M$ FOs, it is possible to generate $N\ll M$ FOs that represent their combined flexibility, with some losses. The opposite process is called disaggregation: from a schedule for an aggregated FOs, it is possible to generate schedules for the original FOs so that their combined energy consumption is the same as the original schedule.
The FO life cycle is shown in Figure 1. Two main actors are involved: an automatic agent representing the prosumer/device owner, which creates and executes the FOs, and another one who manages and optimizes the FOs, which is often either the Balance Responsible Party (BRP) or the aggregator: for clarity, in this paper we will assume it to be the aggregator [20]. FOs are first generated by the prosumer agent from previous forecasts and analysis, and then sent to the aggregator, who decides whether to accept the FO or not and notifies its decision to the prosumer. If the FO is rejected, the process ends here; if it is accepted, the FO is aggregated to other FOs and optimized with them. After this, a schedule is determined for the FO and sent back to the prosumer agent for execution.
Figure 1: A schematic description of the FOs life cycle (taken from [21]).
II-C Standard FlexOffers
The simplest FO constraints are start time constraints and slice (energy) constraints. We are considering time as discrete and divided into time units, also called time slices: their duration is one hour in the running example. Start time constraints define the earliest and latest time at which the load can start. Energy constraints define, at each time unit, the minimum and maximum amount of energy that can be consumed from the considered energy load: in other words, at each time unit $t$, two numbers $emin_{t}$ and $emax_{t}$ are defined such that $emin_{t}\leq e_{t}\leq emax_{t}$, where $e_{t}$ is the energy consumption during time unit $t$. An FO with just these constraints is called a Standard FlexOffer (SFO). We can see an SFO example in Figure 2(a) for the running example for eight time units. The pink bars indicate the amount of energy that can be consumed at each time unit: here, at each time unit, it varies between 0.303 kWh and 0.478 kWh. At each time, the horizontal bars indicate a possible amount of energy consumption $e_{t}$.
[…]
Figure 2: Example SFO (a), TECFO (b), DFO (c) and UFO (d).
II-D Total Energy Constraint FlexOffers
The total energy constraint (TEC) is a more complex type of constraint compared to SFO constraints: it defines a lower ($TE_{min}$) and an upper ($TE_{max}$) bound for the total consumption over all the considered time slices. This can be expressed as $TE_{min}\leq\sum\limits_{t=1}^{T}e_{t}\leq TE_{max}$. An FO with slice and total energy constraints is called a Total Energy Constraint FO (TECFO). Figure 2(b) shows this for the running example: we have the slice constraints of the SFO example, plus a TEC defined by $TE_{min}=2.592$ kWh and $TE_{max}=3.381$ kWh.
II-E Dependency FlexOffers
A dependency energy constraint defines dependencies between time slices. At each time unit $t$, a lower and an upper bound on the amount of consumable energy are defined, depending on how much energy has been consumed until time unit $t-1$. Formally, this can be expressed as $a\cdot(e_{1}+\ldots+e_{t-1})+b\cdot e_{t}\leq c$, where $a,b,c$ are real numbers. An FO with dependency energy constraints is called a Dependency FlexOffer (DFO). We can see an example in Figure 2(c), for the first four time units of the running example: here, at each time unit, the $x$ axis represents the amount of energy consumed up to that time, while the $y$ axis represents the amount of consumable energy at that time. We can see that, in general, the more energy has been consumed up to time $t$, the less energy can be consumed at time $t$, and vice versa.
II-F Uncertain FlexOffers
Uncertain FOs (UFOs) [21] are a type of FOs that consider uncertainty from flexibility devices behavior. Two main types of uncertainty are considered: time and amount uncertainty. At each time unit, time uncertainty refers to the probability for the device to be operative at that time, while amount uncertainty refers to the amount of energy that can be consumed by the device at that time. A UFO is created in two steps. First, uncertainty related to the state (in the running example, temperature) is modeled at each time $t$; second, the probability for each energy value at each time to be feasible is calculated, taking into account both types of uncertainty. Some functions $\{f_{1},\ldots,f_{T}\}$ describing those probabilities will be obtained: those functions will define the UFO. UFOs can be visualized by choosing a probability threshold $P_{0}$. At each time $t$, the energy values having probability at least $P_{0}$ of being available for consumption can be described by intervals. Figure 2(d) shows this visualization for the running example: if we choose $P_{0}=1$, at each time unit, the available energy values are described by the pink bars, while if we choose $P_{0}=0.8$, the available energy values are described by the combined pink and blue bars. This is because at the first time unit, the status of the device is known, and the amount of available flexibility can be determined exactly, while for the next time units, there is more uncertainty (represented here by the blue bars) and the amount of flexibility certainly available (pink bars) is lower.
II-G SAREF and SAREF4ENER
Smart Applications REFerence (SAREF) [^2] https://saref.etsi.org/ [12] is an ontology for IoT devices and solutions published by ETSI [^3] https://www.etsi.org/ in a series of Technical Specifications initially released in 2015 and updated in 2017. Even though its initial objective was to build a reference ontology for appliances relevant to energy efficiency, SAREF can also serve as a reference model to enable better integration of data from various vertical domains in the IoT. SAREF has been extended to different domains: specifically, SAREF for Energy Flexibility (SAREF4ENER) [13] is its extension for the energy domain, and aims to provide a standardized way to describe and interoperate energy-related devices, systems, and services. It allows us to describe devices such as solar panels, electric vehicles, and smart meters, as well as systems like microgrids, and energy management/demand response services. By using SAREF4ENER, developers can create applications that easily integrate and communicate with different energy devices, systems, and services, regardless of the manufacturer or protocol used, which enables more efficient and sustainable energy management.
II-H dCO
The World Wide Web Consortium (W3C) has developed a standard for enabling Internet of Things (IoT) interoperability, whose name is W3C Web of Things (WoT) [22]. In this context, Thing Descriptions (TDs) [23] have been developed: they describe the metadata and interfaces of the considered devices that participate in the WoT. Things are organized in TD Directories (WoT TDDs), which make it possible to register, manage, and search TDs. An important feature for IoT consumers is to send queries to the WoT directory: this is achieved by having a common ontology to annotate TDs; however, before dCO, no ontologies were integrating WoT into smart buildings. dCO [14] has been created with the purpose of achieving this integration: in this paper, we will focus on the aspect of dCO that strictly concerns energy flexibility. A description of the modules can be found in Appendix -A). The contribution of our work is the energy flexibility module, which is not described in [14]: in this paper, we explain how it extends SAREF4ENER.
III dCO Extension of SAREF4ENER
This section shows how SAREF4ENER has been extended to describe FOs. We show how we designed the dCO flexibility module and how we encode attributes and constraints. Details on FO messages can be found in Appendix -B.
III-A FlexOffer Message
As mentioned in Section II, FOs are sent between actors: this is achieved by FO messages, which encode FO metadata and information about energy constraints. The attributes of an FO message describe the FO ID, state (e.g., accepted, rejected), time granularity, geographical location, time limits for acceptance and assignment, constraints on energy, time and price, and an energy schedule. Of these, the only ones to have an equivalent in SAREF4ENER are state, simple energy constraints (slice, TECs), price constraints and duration, while more advanced energy constraints (for DFOs and UFOs) and other attributes had to be added as extensions. Further information can be found in Appendix -B, in Table II for the attributes and Table I for the mapping to SAREF4ENER. In the next section, we show which classes and attributes we created for dCO, in order to model the parts missing in SAREF4ENER.
III-B Energy Flexibility Ontological Design
We now describe how dCO has been designed for flexibility. A complete overview can be seen in Figure 3.
Figure 3: Overview of the flexibility module of dCO.
In the figure, we highlighted in purple the existing SAREF4ENER (written with prefix s4ener) classes that we used, and in yellow the new dCO (written with prefix dco) that we have defined. In this section, we use the notation O:e to describe the element e belonging to the ontology O. The core class for FO representation is dco:FlexOffer: it is defined as a subclass of s4ener:PowerSequence, the SAREF4ENER class describing a power sequence, meaning that dco:FlexOffer inherits the object properties and data properties of s4ener:PowerSequence. Specifically, dco:FlexOffer consists of at least one power sequence slot, represented by s4ener:Slot. Regarding the other dCO classes, dco:FlexOffer consists of at least one profile constraint, denoted by the class dco:FlexOfferProfileConstraint. This class has the following properties:
  • dco:hasEnergyConstraintList: (max 1 s4ener:EnergyMin, max 1 s4ener:EnergyMax). In Figure 2(a), they are 0.303 and 0.478 respectively);
  • s4ene:hasPrice: sarefPrice (contains the list(s) of price constraints for one-time unit, in EUR);
  • s4ener:MinDuration: The minimum duration of a FlexOffer, in time units; in Figure 2(a), it is 8;
  • s4ener:MaxDuration: The maximum duration of a FlexOffer, in time units; in Figure 2(a), it is 8;
  • dco:totalEnergyConstraint: (max 1 s4ener:EnergyMin, max 1 s4ener:EnergyMax). In Figure 2(b), they are 2.592 and 3.381 respectively);
  • dco:dependencyEnergyConstraintList: Contains the dependency constraints for energy in a Dependency FlexOffer;
  • dco:uncertainFunctions: (min 1 Function): Contains the parameters which define the uncertain FlexOffer;
  • dco:uncertainThreshold: a real number that represents the probability threshold. In the running example in Figure 2(d), it can be 1 (pink) or 0.8 (blue).
Also, dco:FlexOffer may have one or more price constraints, denoted by the class dco:FlexOfferPriceConstraint, which has the following properties: dco:Duration: saref:Time; dco:hasPrice: saref:Price.
Finally, dco:FlexOffer can have one or more schedules, represented by dco:Schedule, having the following properties:
  • dco:scheduleType: can be one of the two classes: dco:DefaultSchedule or dco:FlexOfferSchedule;
  • dco:scheduleID: the id of a schedule;
  • dco:updateID: the id of the schedule update;
  • dco:startTime: the start time of a schedule.
Regarding dco:scheduleType, there are two choices because FOs may have two schedules, i.e., a default schedule (dco:defaultSchedule), which indicates a default baseload, or an FO schedule (dco: FlexOfferSchedule), which indicates an FO-optimized schedule. Also, dco:Schedule consists of at least one slice, represented by dco:Slice, which is a subclass of s4ener:Slot. Specifically, dco:Slice inherits all properties of s4ener:Slot, and extends it with the dco:hasPrice object property, which defines the energy price of a slice. In the running example, the SFO shown in Figure 2(a) has eight slices. A dco:slice is associated with the following:
  • s4ener:DefaultDuration: the duration of a dco:Slice;
  • s4ener:hasEnergyValueType: range is either s4ener: Energy or s4ener:Power. Here, we are considering energy;
  • s4ener:hasPrice: the unitary price (in EUR/kWh) of energy in a dco:Slice.
We finally show the properties of the dco:FlexOffer class:
  • dco:hasState: State of the FO (initial/offered/ accepted/rejected/assigned/executed/ invalid/ canceled);
  • dco:stateReason: Reason for FO rejection (if State == rejected);
  • dco:locationID: the location of the FO in the grid system;
  • dco:acceptBeforeInterval: Interval before which FO must be accepted;
  • dco:assignmentBeforeInterval: Interval before which FO must be scheduled;
  • dco:startBeforeInterval / dco:startAfterInterval: Interval before / after which FO must be started;
  • dco:endBeforeInterval / dco:endAfterInterval: Interval before / after which FO execution must end;
  • dco:acceptanceBeforeTime: Absolute time before which FO must be accepted;
  • dco:assignmentBeforeTime: Absolute time before which FO must be scheduled;
  • dco:numSecondsPerInterval: Duration (in seconds) of a time interval. The default value is 900, but in the running example, we consider it as 3600;
  • dco:startBeforeTime / dco:startAfterTime: Absolute time before / after which FO must be started (in the running example, 00:00 of the considered day);
  • dco:endBeforeTime / dco:endAfterTime: Absolute time before / after which FO execution must end (in the running example, 08:00 of the considered day);
  • dco:creationTime: Absolute time at which the FO has been created;
  • dco:priceConstraintStartTime: The start time of the price constraint.
III-C Constraints Encoding
Here, we show how we encode the FO constraints into the dCO ontology: for those constraints, we will use the same notation we used in Section II. The only properties in SAREF4ENER that can be used for this are s4ener:EnergyMin and s4ener:EnergyMax: it is possible to represent SFOs and TECFOs by mapping to them $emin_{t}$ and $emax_{t}$ for slice constraints, and $TE_{min}$ and $TE_{max}$ for TECs. Further information on the FO mapping to SAREF4ENER can be found in Tables III and IV in Appendix -B.
Regarding DFOs, we show how we encoded the constraints. Consider the running example, and the DFO represented in Figure 2(c). The slices are represented by polygons, which are intersections of hyperplanes: thus, we can write these polygons as sets of inequalities over $x$ and $y$. We can then build a matrix, with as many rows as the number of inequalities defining the polygon: in this case, the number of rows would be six. As the inequalities are linear, they can be written as $ax+by\leq c$: the corresponding row of the matrix will then be $[a\;b\;c]$. This is called H-representation of the DFO [24]. We show below the constraints and matrix representation of the fourth slice. The property DependencyEnergyConstraintList encodes the matrix: details are provided in Appendix -E.
$y$ $\geq 0.309$
$y$ $\leq 0.442$
$x$ $\geq 0.972$
$x$ $\leq 1.246$
$0.088\cdot x+y$ $\geq 0.41$
$0.088\cdot x+y$ $\leq 0.537$
$\begin{bmatrix}0&-1&-0.309\\ 0&1&0.442\\ -1&0&-0.972\\ 1&0&1.246\\ -0.088&-1&-0.41\\ 0.088&1&0.537\end{bmatrix}$
Finally, UFOs are described by the functions $\{f_{1},\ldots,f_{T}\}$ mentioned in Section II-F. Describing a generic function is nontrivial; however, for the devices supported by FOs, we noticed that they can be written as $f_{t}=\min\{p_{1},\ldots,p_{N_{t}}\}$, where each $p_{i}$ is a polynomial defined in the range of possible energy values. In the running example, a representation of $f_{1}$ and $f_{2}$ can be seen in Figure 4: those functions are, however, represented on the $y$ axis rather than the $x$ axis, because FOs represent energy over the $y$ axis. We can see that for $t=1$ the function is the constant polynomial $f_{1}=1$, while for $t=2$ we have $f_{2}=\min\{p_{1},p_{2},p_{3}\}$, where $p_{1}=1$, $p_{2}=66.67\cdot x-20.6$, and $p_{3}=-66.67\cdot x+29.467$. Once the polynomials have been determined, they are encoded as a list of lists of coefficients, with the terms corresponding to each degree in increasing order. For example, since $f_{1}=1$, the UFO constraint for $t=1$ is encoded as $[1]$. On the other hand, for $t=2$, the list will be $[[1],[-20.6,66.67],[29.467,-66.67]]$. Details are provided in Appendix -F.
0011110.340.340.340.340.360.360.360.360.380.380.380.38Probabilityt=1Energy,t=1000.50.50.50.511110.350.350.350.350.40.40.40.4Probabilityt=2E,t=2
Figure 4: UFO functions
IV Evaluation
In Section II-G, we mentioned that SAREF4ENER is an ontology that allows for a simple description of energy flexibility. FOs, on the other hand, is a model that represents flexibility in mathematical terms; despite having two different purposes, a parallel can be drawn between SAREF4ENER and FO representations, in that they refer to energy profiles and time slices. However, this parallel highlights a crucial weakness of SAREF4ENER in describing energy flexibility. Specifically, SAREF4ENER is unable to model many necessary attributes for FOs: in the main FO message, only Stateandthe attributes defining slice and total energy constraints can be mapped to SAREF4ENER. It is therefore possible to map SFO and TECFO constraints into SAREF4ENER. However, as mentioned in Section III-A, dependency constraints and UFOs cannot be mapped to SAREF4ENER. On the other hand, the dCO flexibility module has been built in order to provide full support for all types of FOs. This means not only that dCO can represent flexibility through more advanced constraints, but also that it allows for communication via a FO message. This way, by using dCO, developers have the possibility to design applications able to exploit the whole range of benefits coming from FO modeling, i.e., advanced and more accurate flexibility representation, support for more advanced devices, and really fast optimization and (dis)aggregation of flexibility. For example, FOs allow to aggregate, optimize and disaggregate up to $2\cdot 10^{6}$ of loads for 96 time units time horizons [15]. In comparison, an exact approach described in [15] can only optimize up to 500 devices for that time horizon, while other known exact approaches cited in [15] fail already to perform optimization for more than 6 time units.
The domOS project demonstrated dCO feasibility in real-world case studies in Denmark, France, and Switzerland: all functionalities, e.g., querying building descriptions, are operative, and interoperability between smart buildings and services is achieved. The energy services using dCO cut down energy consumption and smoothened load consumption curves for 124 individual customers and 50 days. The cumulative amount of flexibility (load shedding) was several mWh.
Regarding flexibility representation, SAREF4ENER is able at most to accurately represent flexibility for devices that only need this type of constraint. This is the case with wet devices, such as dishwashers and washing machines, as shown in [11]. These devices have a load with a fixed power/amount profile, which is only shifted in time. However, for other devices, SFOs/TECFOs are outclassed by DFOs in terms of accuracy in flexibility representation, as we can see in previous literature. The ability to retain flexibility for a model is calculated by the economic metric [19], defined as the amount of profit that can be obtained by optimizing flexibility with the objective of cost reduction. Specifically, for the charge and discharge process of batteries, DFOs retain $61$% of the flexibility, against the $38$% from TECFOs and $10$% from SFOs [19]. UFOs can perform even better, retaining up to $66.4$% of flexibility, which goes up to $86.8$% of flexibility when imbalance penalties for violating constraints are high (high penalty case) [21]. Regarding EVs, flexibility modeling via DFOs is more accurate than SFOs and TECFOs, as slice constraints and TECs are only a subset of DFO constraints, and therefore the amount of flexibility retained by DFOs is higher: yet, in [21], we see that UFOs can retain $92$% of flexibility, against the $77.3$% retained by DFOs. For the high penalty case, UFOs retain $87.5$% of flexibility, against the $61.8$% from DFOs, which still outperform SFOs and TECFOs. Finally, regarding heat pumps, the comparison between DFOs and SFOs has been done in [24], showing that DFOs can retain more than three times the amount of flexibility from SFOs: in [15] DFOs for heat pumps have been further improved, showing that they can retain up to $98.9$% of flexibility according to the economic metric. The ability for dCO to model the significant amounts of flexibility for these devices ensures that developers can create applications that allow for an accurate description of flexibility for batteries, EVs and heat pumps. Table I summarizes this, showing which use cases are supported by the dCO flexibility module and SAREF4ENER, respectively.
Use case dCO SAREF4ENER
Operating wet devices
Charging and discharging batteries
Charging and discharging EVs
Operating heat pumps
TABLE I: Use cases supported.
Furthermore, FOs are extended over time to represent more use cases and flexibility properties: an example is their capability of modeling different energy vectors from electricity, such as heat [15]. This has been our rationale to extend SAREF4ENER to represent both FOs in general and, more specifically, DFO and UFO constraints: the resulting ontology can represent all the advanced use cases presented in this paper, as well as provide a solid foundation for future ontology extensions based on new types of FOs.
In order to validate dCO, we also used the OntOlogy Pitfall Scanner! (OOPS!) tool [^4] https://oops.linkeddata.es/ to evaluate it. OOPS! [25] is a tool, well-known in the Semantic Web community, designed with the purpose of identifying pitfalls and design errors in ontologies. There are three levels of importance in pitfalls identified by OOPS!. Critical pitfalls need to be resolved, as they may affect the ontology consistency, reasoning, and applicability. Important pitfalls would improve the ontology’s robustness if fixed, but they do not undermine the ontology’s function. Minor pitfalls would improve the ontology’s readability if fixed, but do not pose problems for the ontology. OOPS! identified quite a few minor and important pitfalls, but only 11 critical ones. However, the critical pitfalls were all of the type Defining multiple domains or ranges in properties. In our case, it was not only valid, but also necessary for properties to have multiple domains and ranges to capture the complexity of our flexibility scenario. A simple example of this is the dco:hasEnergyConstraintList property, which can have up to one s4ener:EnergyMin attribute and up to one s4ener:EnergyMax attribute. For this reason, we intentionally designed our ontology with this feature, and therefore none of the critical pitfalls identified by the OOPS! tool undermines the function of the ontology, rather, they are exactly as they should be. Thus, the ontology meets the checked requirements.
Regarding the integration of this ontology with existing standards or protocols, our ontology allows for better flexibility modeling, being able to capture more complex energy constraints and integrate FOs. Our proposed ontology is also compatible with the Common Information Model (CIM) [^5] https://www.entsoe.eu/digital/common-information-model/, which is the standard from the International Electrotechnical Commission for electric power transmission and distribution, and is capable of representing more complex energy constraints than CIM. Finally, FOs integrate with EEBus [^6] https://www.eebus.org/, another energy data standard that builds on SAREF4ENER and only has simple flexibility models.
V Conclusions and Future Work
Energy flexibility models are capable of accurate representations of flexibility, and FOs in particular can do that while being scalable for long time horizons and many devices. For their real-life application, it is important to allow flexibility in data exchange with different types of devices in several contexts, like smart buildings. Standardization of data formats is therefore required, and ontologies fulfill this task. Ontologies have evolved over the years, and SAREF4ENER is the current industry standard ontology for describing the energy usage of smart devices. However, SAREF4ENER fails to support flexibility for relevant use cases such as batteries, EVs, and heat pumps. This paper aims to cover this gap by proposing an extension of SAREF4ENER capable of modeling FOs: we integrated this extension into the dCO ontology, as its flexibility module. This extension allows for the representation of FO message attributes and, specifically, FO constraints: this way, advanced constraints such as dependency and uncertainty constraints can be encoded, and it is possible to accurately represent nontrivial use cases such as the ones mentioned above. Future work will be centered on extending the dCO ontology with newly developed FO features for handling more use cases and flexibility properties. Regarding dCO maintenance, it will follow the evolution of FOs, which evolve but are always backwards-compatible by design. Therefore, maintenance of the dCO flexibility module will only need to encode eventual new energy constraints, while the core part of the message will not need significant changes.
Acknowledgements
This work was supported by the domOS and FEVER projects, funded under the Horizon 2020 programme, GAs 864537 and 894240 respectively, Flexible Energy Denmark, funded by Innovation Fund Denmark (Case No. 8090-00069B), and CLIMETransition, funded under the EU Next Generation programme (CUP F53D23009450001).